Find and simplify a) f(a), b) f(a + h) and f(a+h)-f(a) / h , h≠0 using the given function.

f(x)=x^2-2x+3
1
Expert's answer
2012-12-13T10:34:36-0500
a)f(a)=a^2-2a+3
b)f(a+h)=(a+h)^2-2(a+h)+3=a^2+2ah+h^2-2a-2h+3;
f(a+h)-f(a) / h=a^2+2ah+h^2-2a-2h+3-(a^2-2a+3)/h;